# Wireless Communication for Low-Altitude Economy with UAV Swarm Enabled Two-Level Movable Antenna System

### v1

@@ Abstract
Unmanned aerial vehicle (UAV) is regarded as a key enabling platform for low-altitude economy, due to its advantages such as three-dimensional (3D) maneuverability, flexible deployment, and line-of-sight (LoS) air-to-air/ground communication links. In particular, the intrinsic high mobility renders UAV especially suitable for operating as a movable antenna (MA) from the sky. In this paper, by exploiting the flexible mobility of UAV swarm and antenna position adjustment of MA, we propose a novel UAV swarm enabled two-level MA system, where UAVs not only individually deploy a local MA array, but also form a larger-scale MA system with their individual MA arrays via swarm coordination. We formulate a general optimization problem to maximize the minimum achievable rate over all ground user equipments (UEs), by jointly optimizing the 3D UAV swarm placement positions, their individual MAs’ positions (or local positions), and receive beamforming for different UEs. To gain useful insights, we first consider the special case where each UAV has only one antenna, under different scenarios of one single UE, two UEs, and arbitrary number of UEs. In particular, for the two-UE case, we derive the optimal UAV swarm placement positions in closed-form that achieves inter-UE interference (IUI)-free communication, where the UAV swarm forms a uniform sparse array (USA) satisfying collision avoidance constraint. While for the general case with arbitrary number of UEs, we propose an efficient alternating optimization algorithm to solve the formulated non-convex optimization problem. Then, we extend the results to the case where each UAV is equipped with multiple antennas. It is shown that an architecture resembling a modular array can achieve the IUI-free communication in the two-UE case. Numerical results verify that the proposed low-altitude UAV swarm enabled MA system significantly outperforms various benchmark schemes, thanks to the exploitation of two-level mobility to create more favorable channel conditions for multi-UE communications.
@@ I Introduction
Note that compared to MAs, UAVs possess even more flexible and wider-range movement, and thus an interesting new idea is to utilize UAV swarm to form MA array. Moreover, different from existing research on MA that only optimizes antenna positions, UAV swarm enabled MA can fully unlock its mobility advantage, where the array geometry is able to be dynamically reconfigured via trajectory optimization, thus yielding a varying array geometry. Preliminary efforts have been devoted to the single UAV-mounted MA [37, 38, 39, 40], and single UAV-mounted passive 6DMA [41], i.e., by carrying an intelligent reflecting surface (IRS) on a UAV [42]. Nevertheless, these works neither take into account the UAV swarm scenario, nor explore the capability of MA system enabled by UAV swarm.
@@
In this paper, motivated by the inherent mobility of UAV swarm and antenna position adjustment of MA, we propose a novel UAV swarm enabled two-level MA system. Specifically, by individually deploying a local MA array on each UAV, all the spatially distributed UAVs cooperatively construct a larger-scale MA system according to their topology. Compared to existing MA systems, UAV swarm enabled MA system involves the dual-scale antenna spatial movement, i.e., the small-scale antenna movement within each local MA array, e.g., at wavelength level, and the large-scale UAV movement, so as to better harness the spatial variations of wireless channels in different scales. By considering low-altitude UAV swarm enabled uplink multi-UE communication, we aim to maximize the minimum achievable rate over all ground UEs. The main contributions of this paper are summarized as follows.
@@
  • First, we propose a novel low-altitude UAV swarm enabled two-level MA system, where each UE channel depends on both UAV swarm placement positions and their mounted MA array positions. Subject to the collision avoidance constraint for UAVs [3], as well as mutual coupling avoidance constraint for UAV-mounted MA array, we formulate an optimization problem to maximize the minimum achievable rate over all ground UEs, by jointly optimizing the 3D UAV swarm placement positions, their individual MAs’ positions, and receive beamforming for all UEs.
@@
  • Second, to gain useful insights, we consider the special case of single-antenna UAV. It is shown that for the single UE communication, the resulting signal-to-noise ratio (SNR) after applying the optimal maximal-ratio combining (MRC) beamforming only depends on the number of UAVs, irrespective of their array geometry. Then, for the two-UE communication, we derive the optimal UAV swarm placement positions that achieves the maximum communication rate without suffering any inter-UE interference (IUI). The result shows that under the collision avoidance constraint, UAV swarm forms a uniform sparse array (USA) [43] for the IUI-free communication. Moreover, for arbitrary number of UEs, we propose an efficient alternating optimization algorithm to solve the highly non-convex problem, where the optimal receive beamforming is derived in closed-form for given UAV swarm placement positions, and for given receive beamforming, the UAV swarm placement optimization problem can be solved via the successive convex approximation (SCA) technique [3].
@@
  • Last, we consider the general case of multi-antenna UAV, where each antenna can move independently within the specified region at its mounted UAV. For the single and two-UE communications, similar results can be obtained as the scenario of single-antenna UAV. In particular, we also consider the special case where all the UAVs adopt the USA [44], the IUI-free communication can be achieved in the two-UE case via joint UAV placement positions and sparsity level adjustment, and the UAV swarm forms an architecture resembling a modular array [45]. Next, the alternating optimization method is applied to solve the problem with arbitrary number of UEs. Numerical results are presented to validate the superior performance of the proposed UAV swarm enabled MA system over various benchmarks, thanks to the two-level mobility to create more favorable channel conditions.
@@
Notations: Scalars are denoted by italic letters. Vectors and matrices are denoted by bold-face lower- and upper-case letters, respectively. The spaces of $M\times N$ complex-valued matrices are denoted as ${{\mathbb{C}}^{M\times N}}$. For a vector ${\bf{x}}$, $\left\|{\bf{x}}\right\|$ denotes its Euclidean norm. For an arbitrary-size matrix ${\bf A}$, its complex conjugate, transpose, and Hermitian transpose are denoted by ${\bf A}^{*}$, ${\bf A}^{T}$, ${\bf A}^{H}$, respectively. The spectral norm and Frobenius norm of ${\bf A}$ are denoted as ${\left\|{\bf{A}}\right\|_{\_}2}$ and ${\left\|{\bf{A}}\right\|_{\_}F}$, respectively. The distribution of a circularly symmetric complex Gaussian random vector with mean $\bf{x}$ and covariance matrix $\bf{\Sigma}$ is denoted as ${\cal CN}\left({\bf{x},\bf{{\Sigma}}}\right)$, and $\sim$ stands for “distributed as”. For a complex-valued number $x$, ${\mathop{\rm Re}\nolimits}\left\{x\right\}$ represents its real part. $\bmod\left({a,b}\right)$ returns the remainder after division of $a$ by $b$. The symbol ${\rm j}$ denotes the imaginary unit of complex numbers, with ${{\rm j}^{2}}=-1$. ${\cal O}\left({\cdot}\right)$ denotes the standard big-O notation.
@@ II System Model And Problem Formulation
As illustrated in Fig. 1, we consider a low-altitude UAV swarm enabled wireless communication system, where $L$ UAVs cooperatively serve $K$ single-antenna ground UEs. Each UAV is equipped with a linear MA array with $M$ antenna elements, while a flexible two-level MA system is formed by dynamically moving each UAV and the positions of its individual MAs. Without loss of generality, we consider a 3D Cartesian coordinate system, where the location of UE $k$ is denoted as ${\bf w}_{\_}k=\left[{{x_{\_}k},{y_{\_}k},0}\right]^{T}$, $k\in{\cal K}$, with ${\cal K}\triangleq\left[1,\cdots,K\right]$. Let the first antenna be the reference point for each UAV, with its coordinate for UAV $l$ denoted as ${{\bf{q}}_{\_}l}={\left[{{x_{\_}l},{y_{\_}l},{z_{\_}l}}\right]^{T}}$, $1\leq l\leq L$. The coordinate of antenna $m$ for UAV $l$ is ${{\bf{q}}_{\_}{l,m}}={{\bf{q}}_{\_}l}+{{\bf{b}}_{\_}{l,m}}$, where ${{\bf{b}}_{\_}{l,m}}\triangleq{\left[{{{b_{\_}{l,m}}},0,0}\right]^{T}}$ and we assume ${b_{\_}{l,1}}=0$, $\forall l$, without loss of generality. As such, position of antenna $m$ can be dynamically changed by adjusting ${b_{\_}{l,m}}$ for UAV $l$. This thus enables a two-level MA system, realized by the joint control of UAVs’ movement and their individual antennas’ movement.
In practice, the collision avoidance constraints need to be satisfied by UAVs, which is given by [3]
$\left\|{{{\bf{q}}_{\_}l}-{{\bf{q}}_{\_}{l^{\prime}}}}\right\|\geq{d_{\_}{\min}},\ \forall l,l^{\prime}\neq l,$ (1)
where $d_{\_}{\min}$ denotes the minimum distance to avoid the collision between UAVs. Besides, to avoid the mutual coupling between adjacent MA elements at each UAV, the minimum distance $_{\tilde{d}}_{\_}{\min}$ is required, which yields the following constraints,
$\left\|{{{\bf{q}}_{\_}{l,m}}-{{\bf{q}}_{\_}{l,m^{\prime}}}}\right\|=\left|{{b_{\_}{l,m}}-{b_{\_}{l,m^{\prime}}}}\right|\geq{{\tilde{d}}_{\_}{\min}},\forall l,m,m^{\prime}\neq m,$ (2)
where $_{\tilde{d}}_{\_}{\min}$ is in practice much smaller than $d_{\_}{\min}$.
@@
Let ${{\bf{q}}_{\_}r}={\left[{{x_{\_}r},{y_{\_}r},{z_{\_}r}}\right]^{T}}$ denote the reference location for the $L$ UAVs. We consider the far-field uniform plane wave (UPW) model for the channels between UAVs and ground UEs. Moreover, let ${{\bm{\kappa}}_{\_}k}={\left[{{\Phi_{\_}k},{\Psi_{\_}k},{\Theta_{\_}k}}\right]^{T}}$ denote the wave propagation direction from ${\bf q}_{\_}r$ to UE $k$, with ${\Phi_{\_}k}\triangleq\sin{\theta_{\_}k}\cos{\phi_{\_}k}$, ${\Psi_{\_}k}\triangleq\sin{\theta_{\_}k}\sin{\phi_{\_}k}$, and ${\Theta_{\_}k}\triangleq\cos{\theta_{\_}k}$, respectively, where ${{\theta_{\_}k}}$ and ${{\phi_{\_}k}}$ denote the elevation and azimuth angles of arrival (AoAs), respectively, as illustrated in Fig. 1. The difference of the wavefront propagation distance between ${\bf q}_{\_}r$ and ${{\bf{q}}_{\_}{l,m}}$ for UE $k$ is
${\varphi_{\_}k}\left({{{\bf{q}}_{\_}{l,m}}}\right)$ $={\bm{\kappa}}_{\_}k^{T}\left({{{\bf{q}}_{\_}{l,m}}-{{\bf{q}}_{\_}r}}\right)=\left({{x_{\_}l}+{b_{\_}{l,m}}}\right){\Phi_{\_}k}$ (3)
$\ \ \ +{y_{\_}l}{\Psi_{\_}k}+{z_{\_}l}{\Theta_{\_}k}-{\bm{\kappa}}_{\_}k^{T}{{\bf{q}}_{\_}r}.$
Note that the communication links between UAVs and ground UEs are of high probability to be over LoS channels in practice. The receive array response vector of UAV $l$ for UE $k$ is given by
${{\bf{a}}_{\_}{k,l}}\left({\left\{{{{\bf{q}}_{\_}{l,m}},\forall m}\right\}}\right)={\left[{{e^{{\rm{j}}\frac{{2\pi}}{\lambda}{\varphi_{\_}k}\left({{{\bf{q}}_{\_}{l,1}}}\right)}},\cdots,{e^{{\rm{j}}\frac{{2\pi}}{\lambda}{\varphi_{\_}k}\left({{{\bf{q}}_{\_}{l,M}}}\right)}}}\right]^{T}}.$ (4)
@@
Moreover, the $L$ UAVs cooperate to form an $LM$-element distributed antenna array, and thus the channel from the UAV swarm enabled MA system to UE $k$ is
${{\bf{h}}_{\_}k}\left({\left\{{{{\bf{q}}_{\_}{l,m}},\forall l,m}\right\}}\right)={\alpha_{\_}k}{{\bf{a}}_{\_}k}\left({\left\{{{{\bf{q}}_{\_}{l,m}},\forall l,m}\right\}}\right),$ (5)
@@
where ${\alpha_{\_}k}=\sqrt{{\beta_{\_}0}}/{d_{\_}{r,k}}$ denotes the channel gain of UE $k$, with ${\beta_{\_}0}$ being the channel power at the reference distance of ${d_{\_}0}=1$ m and ${d_{\_}{r,k}}=\left\|{{{\bf{w}}_{\_}k}-{{\bf{q}}_{\_}r}}\right\|$ being the distance between ${\bf q}_{\_}r$ and ${\bf w}_{\_}k$, respectively. Besides, ${{\bf{a}}_{\_}k}\left({\left\{{{{\bf{q}}_{\_}{l,m}},\forall l,m}\right\}}\right)\in{\mathbb{C}}^{LM\times 1}$ denotes the receive array response vector of UE $k$, given by
${{\bf{a}}_{\_}k}\left({\left\{{{{\bf{q}}_{\_}{l,m}},\forall l,m}\right\}}\right)=\left[{{\bf{a}}_{\_}{k,1}^{T}\left({\left\{{{{\bf{q}}_{\_}{1,m}},\forall m}\right\}}\right),\cdots,}\right.$ (6)
${\left.{{\bf{a}}_{\_}{k,L}^{T}\left({\left\{{{{\bf{q}}_{\_}{L,m}},\forall m}\right\}}\right)}\right]^{T}}.$
@@
In the following, for brevity, we use the notations ${{\bf{h}}_{\_}k}$, ${{\bf{a}}_{\_}k}$ and ${{\bf{a}}_{\_}{k,l}}$ to replace ${{\bf{h}}_{\_}k}\left({\left\{{{{\bf{q}}_{\_}{l,m}},\forall l,m}\right\}}\right)$, ${{\bf{a}}_{\_}k}\left({\left\{{{{\bf{q}}_{\_}{l,m}},\forall l,m}\right\}}\right)$ and ${{\bf{a}}_{\_}{k,l}}\left({\left\{{{{\bf{q}}_{\_}{l,m}},\forall m}\right\}}\right)$, respectively.
@@
Denote by ${s_{\_}k}$ the independent and identically distributed (i.i.d.) information-bearing symbol of UE $k$, with ${\mathbb{E}}[{{{\left|{{s_{\_}k}}\right|}^{2}}}]=1$. To detect its symbol, the receive beamforming ${{\bf{v}}_{\_}k}\in{\mathbb{C}}^{LM\times 1}$ is applied at the UAV swarm enabled MA system, with $\left\|{{\bf{v}}_{\_}k}\right\|=1$. Thus, the resulting signal of UE $k$ after applying the receive beamforming is
${y_{\_}k}={\bf{v}}_{\_}k^{H}{{\bf{h}}_{\_}k}\sqrt{{P_{\_}k}}{s_{\_}k}+{\bf{v}}_{\_}k^{H}\sum\limits_{\_}{i=1,i\neq k}^{K}{{{\bf{h}}_{\_}i}\sqrt{{P_{\_}i}}{s_{\_}i}}+{\bf{v}}_{\_}k^{H}{\bf{z}},$ (7)
@@
where ${{P_{\_}i}}$, $i=1,\cdots,K$, denotes the transmit power of UE $i$, ${\bf{z}}\sim{\cal CN}\left({{\bf{0}},{\sigma^{2}}{{\bf{I}}_{\_}{LM}}}\right)$, with each element denoting the i.i.d. additive white Gaussian noise (AWGN) with zero mean and power ${\sigma^{2}}$.

### v2

@@ Abstract
Unmanned aerial vehicle (UAV) is regarded as a key enabling platform for low-altitude economy, due to its advantages such as three-dimensional (3D) maneuverability, flexible deployment, and line-of-sight (LoS) air-to-air/ground communication links. In particular, the intrinsic high mobility renders UAV especially suitable for operating as a movable antenna (MA) from the sky. In this paper, by exploiting the flexible mobility of UAV swarm and antenna position adjustment of MA, we propose a novel UAV swarm enabled two-level MA system, where UAVs not only individually deploy a local MA array, but also form a larger-scale MA system with their individual MA arrays via swarm coordination. We formulate a general optimization problem to maximize the minimum achievable rate over all ground user equipments (UEs), by jointly optimizing the 3D UAV swarm placement positions, their individual MAs’ positions (or local positions), and receive beamforming for different UEs. To gain useful insights, we first consider the special case where each UAV has only one antenna, under different scenarios of one single UE, two UEs, and arbitrary number of UEs. In particular, for the two-UE case, we derive the optimal UAV swarm placement positions in closed-form that achieves inter-UE interference (IUI)-free communication when the uniform plane wave (UPW) model holds, where the UAV swarm forms a uniform sparse array (USA) satisfying minimum safe distance constraint. While for the general case with arbitrary number of UEs, we propose an efficient alternating optimization algorithm to solve the formulated non-convex optimization problem. Then, we extend the results to the case where each UAV is equipped with multiple antennas. Numerical results verify that the proposed low-altitude UAV swarm enabled MA system significantly outperforms various benchmark schemes, thanks to the exploitation of two-level mobility to create more favorable channel conditions for multi-UE communications.
@@ I Introduction
Note that compared to MAs, UAVs possess even more flexible and wider-range movement, and thus an interesting new idea is to utilize UAV swarm to form MA array. Moreover, different from existing research on MA that only optimizes antenna positions, UAV swarm enabled MA can fully unlock its mobility advantage, where the array geometry is able to be dynamically reconfigured via trajectory optimization, thus yielding a varying array geometry. Preliminary efforts have been devoted to the single UAV-mounted MA [37, 38, 39, 40]. Specifically, a directional MA was mounted on a UAV to minimize the total data collection time in a backscatter sensor network in [37]. In [38] and [39], the authors utilized the single UAV-mounted MA array to facilitate the minimum beamforming gain for secondary users (SUs) and enhance the sum rate of all UEs, respectively. In [40], the 6DMA was deployed at the single UAV to mitigate the interference caused by the co-channel terrestrial transmissions. Besides the active antenna, the authors in [41] utilized the single UAV-mounted passive 6DMA, i.e., by carrying an intelligent reflecting surface (IRS) on a UAV [42], to maximize the minimum received signal-to-noise ratio (SNR) among all UEs. Nevertheless, these works neither take into account the UAV swarm scenario, nor explore the capability of MA system enabled by UAV swarm. Meanwhile, it is also worth pointing out that the UAV swarm enabled MA system may face several practical challenges, such as the limited endurance of UAV swarm due to the finite onboard battery, positional inaccuracy, array misalignment, orientation jitter, and synchronization issues, which deserve further investigations.
@@
In this paper, motivated by the inherent mobility of UAV swarm and antenna position adjustment of MA, we propose a novel UAV swarm enabled two-level MA system. Specifically, by individually deploying a local MA array on each UAV, all the spatially distributed UAVs cooperatively construct a larger-scale MA system according to their topology. Compared to existing MA systems, UAV swarm enabled MA system involves the dual-scale antenna spatial movement, i.e., the small-scale antenna movement within each local MA array, e.g., at wavelength level, and the large-scale UAV movement, so as to better harness the spatial variations of wireless channels in different scales. In particular, the UAV swarm enabled MA system can be deployed on demand, which is appealing for assorted application scenarios, such as information dissemination and data collection, data offloading in hotspot areas, and rapid communication restoration after infrastructure failure [3]. Moreover, benefiting from the desired LoS air-to-air/ground links, UAV swarm enabled MA system can provide the ISAC service efficiently, so as to support low-altitude economy applications such as logistics and delivery as well as environmental monitoring. In this paper, by considering low-altitude UAV swarm enabled uplink multi-UE communication, we aim to maximize the minimum achievable rate over all ground UEs. The main contributions of this paper are summarized as follows.
@@
  • First, we propose a novel low-altitude UAV swarm enabled two-level MA system, where each UE channel depends on both UAV swarm placement positions and their mounted MA array positions. Subject to the minimum safe distance constraint for UAVs [3], as well as mutual coupling avoidance constraint for UAV-mounted MA array, we formulate an optimization problem to maximize the minimum achievable rate over all ground UEs, by jointly optimizing the 3D UAV swarm placement positions, their individual MAs’ positions, and receive beamforming for all UEs.
@@
  • Second, to gain useful insights, we consider the special case of single-antenna UAV. It is shown that for the single UE communication, the resulting SNR after applying the optimal maximal-ratio combining (MRC) beamforming only depends on the number of UAVs, irrespective of their array geometry. Then, for the two-UE communication, we derive the optimal UAV swarm placement positions that achieves the maximum communication rate without suffering any inter-UE interference (IUI) when the uniform plane wave (UPW) model holds. The result shows that under the minimum safe distance constraint, UAV swarm forms a uniform sparse array (USA) [43] for the IUI-free communication. Moreover, for arbitrary number of UEs, we propose an efficient alternating optimization algorithm to solve the highly non-convex problem, where the optimal receive beamforming is derived in closed-form for given UAV swarm placement positions, and for given receive beamforming, the UAV swarm placement optimization problem can be solved via the successive convex approximation (SCA) technique [3].
@@
  • Last, we consider the general case of multi-antenna UAV, where each antenna can move independently within the specified region at its mounted UAV. For the single and two-UE communications, similar results can be obtained as the scenario of single-antenna UAV. In particular, we also consider the special case where all the UAVs adopt the USA [44], the IUI-free communication can be achieved in the two-UE case via joint UAV placement positions and sparsity level adjustment. Next, the alternating optimization method is applied to solve the problem with arbitrary number of UEs. Numerical results are presented to validate the superior performance of the proposed UAV swarm enabled MA system over various benchmarks, thanks to the two-level mobility to create more favorable channel conditions.
@@
Notations: Scalars are denoted by italic letters. Vectors and matrices are denoted by bold-face lower- and upper-case letters, respectively. The spaces of $M\times N$ complex-valued matrices are denoted as ${\mathbb{C}}^{M\times N}$}. For a vector $\bf{x}$}, $\left\|{\bf{x}}\right\|$ denotes its Euclidean norm. For an arbitrary-size matrix $\bf A$}, its complex conjugate, transpose, and Hermitian transpose are denoted by $\bf A$$^{*}$, ${\bf A}^{T}$, ${\bf A}^{H}$, respectively. The spectral norm and Frobenius norm of $\bf A$} are denoted as $\left\|{\bf{A}}\right\|_{2}$} and $\left\|{\bf{A}}\right\|_{F}$}, respectively. The distribution of a circularly symmetric complex Gaussian random vector with mean $\bf{x}$ and covariance matrix $\bf{\Sigma}$ is denoted as ${\cal CN}\left({\bf{x},\bf{{\Sigma}}}\right)$, and $\sim$ stands for “distributed as”. For a complex-valued number $x$, ${\mathop{\rm Re}\nolimits}\left\{x\right\}$ represents its real part. $\bmod\left({a,b}\right)$ represents the integer modulo operation, which returns the remainder after division of $a$ by $b$. The symbol $\rm j$} denotes the imaginary unit of complex numbers, with ${{\rm j}^{2}}=-1$. ${\cal O}\left({\cdot}\right)$ denotes the standard big-O notation.
@@ II System Model And Problem Formulation
As illustrated in Fig. 1, we consider a low-altitude UAV swarm enabled wireless communication system, where $L$ UAVs cooperatively serve $K$ single-antenna ground UEs. Each UAV is equipped with a linear MA array with $M$ antenna elements, while a flexible two-level MA system is formed by dynamically moving each UAV and the positions of its individual MAs. Without loss of generality, we consider a 3D Cartesian coordinate system, where the location of UE $k$ is denoted as ${\bf w}_{k}=\left[{{x_{k}},{y_{k}},0}\right]^{T}$, $k\in{\cal K}$, with ${\cal K}\triangleq\left[1,\cdots,K\right]$. Let the first antenna be the reference point for each UAV, with the coordinate of UAV $l$ denoted as ${{{\bf{\bar{q}}}}_{l}}={\left[{{{\bar{x}}_{l}},{{\bar{y}}_{l}},{{\bar{z}}_{l}}}\right]^{T}}$, $1\leq l\leq L$. Moreover, the coordinate of antenna $m$ for UAV $l$ is ${{\bf{\bar{q}}}_{l,m}}={{\bf{\bar{q}}}_{l}}+{{\bf{b}}_{l,m}}$, where ${{\bf{b}}_{l,m}}\triangleq{\left[{{{b_{l,m}}},0,0}\right]^{T}}$ and we assume ${b_{l,1}}=0$, $\forall l$, without loss of generality. As such, position of antenna $m$ can be dynamically changed by adjusting $b_{l,m}$} for UAV $l$. This thus enables a two-level MA system, realized by the joint control of UAVs’ movement and their individual antennas’ movement.
@@
Let ${{\bf{q}}_{r}}={\left[{{x_{r}},{y_{r}},{z_{r}}}\right]^{T}}$ denote the reference location for the $L$ UAVs. Let ${{\bm{\kappa}}_{k}}={\left[{{\Phi_{k}},{\Psi_{k}},{\Theta_{k}}}\right]^{T}}$ denote the wave propagation direction from UE $k$ to ${\bf q}_{r}$, with ${\Phi_{k}}\triangleq\sin{\theta_{k}}\cos{\phi_{k}}$, ${\Psi_{k}}\triangleq\sin{\theta_{k}}\sin{\phi_{k}}$, and ${\Theta_{k}}\triangleq\cos{\theta_{k}}$, respectively, where $\theta_{k}$}} and $\phi_{k}$}} denote the elevation and azimuth angles of arrival (AoAs), respectively, as illustrated in Fig. 1. Moreover, let ${{\bf{q}}_{l}}={\left[{{x_{l}},{y_{l}},{z_{l}}}\right]^{T}}$ denote the position of UAV $l$ relative to the reference point ${\bf{q}}_{r}$}, and it follows that ${{{\bf{\bar{q}}}}_{l}}={{\bf{q}}_{r}}+{{\bf{q}}_{l}}$. Similarly, we have ${{{\bf{\bar{q}}}}_{l,m}}={{\bf{q}}_{r}}+{{\bf{q}}_{l,m}}$, with ${{\bf{q}}_{l,m}}={{\bf{q}}_{l}}+{{\bf{b}}_{l,m}}$ denoting the position of antenna $m$ for UAV $l$ relative to the reference point ${\bf{q}}_{r}$}. We assume that each UAV knows its relative position to the reference point, which can be obtained via satellite-based localization systems, e.g., global positioning system (GPS) or Beidou.
In practice, the minimum safe distance constraints need to be satisfied by UAVs, which is given by [3]
$\left\|{{{\bf{q}}_{l}}-{{\bf{q}}_{l^{\prime}}}}\right\|\geq{d_{\min}},\ \forall l,l^{\prime}\neq l,$ (1)
where $d_{\min}$ denotes the minimum distance to guarantee the safe operation between UAVs. Besides, to avoid the mutual coupling between adjacent MA elements at each UAV, the minimum distance ${\tilde{d}}_{\min}$ is required, which yields the following constraints,
$\left\|{{{\bf{q}}_{l,m}}-{{\bf{q}}_{l,m^{\prime}}}}\right\|=\left|{{b_{l,m}}-{b_{l,m^{\prime}}}}\right|\geq{{\tilde{d}}_{\min}},\forall l,m,m^{\prime}\neq m,$ (2)
where ${\tilde{d}}_{\min}$ is in practice much smaller than $d_{\min}$.
Note that the communication links between UAVs and ground UEs are of high probability to be over LoS channels in practice. We consider the general near-field spherical wavefront model for the channels between UAVs and ground UEs, which includes the far-field UPW model as a special case. Meanwhile, the signals from UE to the local MA array of each UAV can be well approximated as UPW, since the movable region of the local MA array is much smaller than the link distance. Let ${{\bm{\kappa}}_{k,l}}={\left[{{\Phi_{k,l}},{\Psi_{k,l}},{\Theta_{k,l}}}\right]^{T}}$ denote the wave propagation direction from UE $k$ to ${\bf{\bar{q}}}_{l}$}, where ${\Phi_{k,l}}=\sin{\theta_{k,l}}\cos{\phi_{k,l}}$, ${\Psi_{k,l}}=\sin{\theta_{k,l}}\sin{\phi_{k,l}}$, and ${\Theta_{k,l}}=\cos{\theta_{k,l}}$, with $\theta_{k,l}$} and $\phi_{k,l}$} being the elevation and azimuth AoAs, respectively. For UAV $l$, the difference of the wavefront propagation distance between ${\bf{\bar{q}}}_{l}$} and ${\bf{\bar{q}}}_{l,m}$} for UE $k$ is
${\varphi_{k}}\left({{{{\bf{\bar{q}}}}_{l,m}}}\right)={\bm{\kappa}}_{k,l}^{T}\left({{{{\bf{\bar{q}}}}_{l,m}}-{{{\bf{\bar{q}}}}_{l}}}\right)={\bm{\kappa}}_{k,l}^{T}{{\bf{b}}_{l,m}}={\Phi_{k,l}}{b_{l,m}},$ (3)
which is the projection length of the vector $\left({{{{\bf{\bar{q}}}}_{l,m}}-{{{\bf{\bar{q}}}}_{l}}}\right)$ onto the wave propagation direction vector ${\bm{\kappa}}_{k,l}$}. Then, the receive array response vector of UAV $l$ for UE $k$ is given by [35]
${{\bf{a}}_{k,l}}\left({\left\{{{{{\bf{\bar{q}}}}_{l,m}},\forall m}\right\}}\right)={e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{r_{k,l}}}}\times$ (4)
$\ \ \ \ \ \ {\left[{{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{\varphi_{k}}\left({{{{\bf{\bar{q}}}}_{l,1}}}\right)}},\cdots,{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{\varphi_{k}}\left({{{{\bf{\bar{q}}}}_{l,M}}}\right)}}}\right]^{T}},$
where ${r_{k,l}}=\left\|{{{\bf{\bar{q}}}_{l}}-{{\bf{w}}_{k}}}\right\|$ denotes the link distance between ${\bf{\bar{q}}}_{l}$}} and ${\bf{w}}_{k}$}}. By applying the second-order Taylor approximation, we have
${r_{k,l}}$ $\approx{r_{k}}+{x_{l}}{\Phi_{k}}+{y_{l}}{\Psi_{k}}+{z_{l}}{\Theta_{k}}+$ (5)
$\ \ \ \frac{{x_{l}^{2}+y_{l}^{2}+z_{l}^{2}-{{\left({{x_{l}}{\Phi_{k}}+{y_{l}}{\Psi_{k}}+{z_{l}}{\Theta_{k}}}\right)}^{2}}}}{{2{r_{k}}}},$
where ${r_{k}}=\left\|{{{\bf{w}}_{k}}-{{\bf{q}}_{r}}}\right\|$ denotes the distance between ${\bf q}_{r}$ and ${\bf w}_{k}$.
@@
Moreover, the $L$ UAVs cooperate to form an $LM$-element distributed antenna array, and thus the channel from UE $k$ to the UAV swarm enabled MA system is
${{\bf{h}}_{k}}\left({\left\{{{{\bf{\bar{q}}}_{l,m}},\forall l,m}\right\}}\right)={\alpha_{k}}{{\bf{a}}_{k}}\left({\left\{{{{\bf{\bar{q}}}_{l,m}},\forall l,m}\right\}}\right),$ (6)
@@
where ${\alpha_{k}}=\sqrt{{\beta_{0}}}/{r_{k}}$ denotes the channel gain of UE $k$, with $\beta_{0}$} being the channel power at the reference distance of ${d_{0}}=1$ m. Besides, ${{\bf{a}}_{k}}\left({\left\{{{{\bf{\bar{q}}}_{l,m}},\forall l,m}\right\}}\right)\in{\mathbb{C}}^{LM\times 1}$ denotes the receive array response vector of UE $k$, given by
${{\bf{a}}_{k}}\left({\left\{{{{\bf{\bar{q}}}_{l,m}},\forall l,m}\right\}}\right)=\left[{{\bf{a}}_{k,1}^{T}\left({\left\{{{{\bf{\bar{q}}}_{1,m}},\forall m}\right\}}\right),\cdots,}\right$. (7)
${\left.{{\bf{a}}_{k,L}^{T}\left({\left\{{{{\bf{\bar{q}}}_{L,m}},\forall m}\right\}}\right)}\right]^{T}}.$
@@
In the following, for brevity, we use the notations ${\bf{h}}_{k}$}, ${\bf{a}}_{k}$} and ${\bf{a}}_{k,l}$} to replace ${{\bf{h}}_{k}}\left({\left\{{{{\bf{\bar{q}}}_{l,m}},\forall l,m}\right\}}\right)$, ${{\bf{a}}_{k}}\left({\left\{{{{\bf{\bar{q}}}_{l,m}},\forall l,m}\right\}}\right)$ and ${{\bf{a}}_{k,l}}\left({\left\{{{{\bf{\bar{q}}}_{l,m}},\forall m}\right\}}\right)$, respectively.
@@
Denote by $s_{k}$} the independent and identically distributed (i.i.d.) information-bearing symbol of UE $k$, with ${\mathbb{E}}[{{{\left|{{s_{k}}}\right|}^{2}}}]=1$. To detect its symbol, the receive beamforming ${{\bf{v}}_{k}}\in{\mathbb{C}}^{LM\times 1}$ is applied at the UAV swarm enabled MA system, with $\left\|{{\bf{v}}_{k}}\right\|=1$. Thus, the resulting signal of UE $k$ after applying the receive beamforming is
${y_{k}}={\bf{v}}_{k}^{H}{{\bf{h}}_{k}}\sqrt{{P_{k}}}{s_{k}}+{\bf{v}}_{k}^{H}\sum\limits_{i=1,i\neq k}^{K}{{{\bf{h}}_{i}}\sqrt{{P_{i}}}{s_{i}}}+{\bf{v}}_{k}^{H}{\bf{z}},$ (8)
@@
where $P_{i}$}}, $i=1,\cdots,K$, denotes the transmit power of UE $i$, ${\bf{z}}\sim{\cal CN}\left({{\bf{0}},{\sigma^{2}}{{\bf{I}}_{LM}}}\right)$, with each element denoting the i.i.d. additive white Gaussian noise (AWGN) with zero mean and power $\sigma^{2}$}. To reveal the fundamental performance limit, synchronization is assumed for the UAV swarm system, which can be achieved via a synchronization protocol as in [45].